A cube has surface area 6x^2 + 12x +6. Find the
length of a side.

Answer by jim_thompson5910(35256)

6x^2 + 12x +6 = 6(x^2+2x+1) = 6(x+1)^2

Since the surface area of a cube with side 's' is 6s^2, this means that the side length is x+1 units.